a square with one side that measures 5 1/2 inches. I don't get the whole P=2(lxw) thing....please help!
Answer by stanbon(48546)

You have the formula for the Perimeter (distance around) of a rectangle.
P=2l +2w or P=2(l+w)
Since you have a square the length and the width are the same.
The perimeter in your case is p=2(5 1/2)+2(5 1/2)
P=11+11= 22
Cheers,
stan H.